# Adiabatic Electron Transfer in the Barrierless and Marcus-Inverted Regimes

### v4

@@ Abstract
Here it is shown that in the adiabatic limit of condensed-phase electron transfer, the onset of barrierless transition occurs at a lower driving force than predicted by the non-adiabatic Marcus formulation. Furthermore, in the adiabatic limit of the Marcus-inverted region, the standard mechanism of electron transfer becomes topologically forbidden. This behavior arises from a topological change in the mapping between the adiabatic and diabatic electronic surfaces, emerging precisely at the onset of the Marcus-inverted region. In this case, alternative mechanisms such as tunneling and non-radiative decay may dominate the rate, typically orders of magnitude slower than the rate calculated from Marcus theory.

### v5

@@ Abstract
While the Marcus theory of electron transfer defines the reorganization energy as a classical quantity, recent work has shown for the first time that the reorganization energy is actually a quantum mechanical object that depends on the electronic coupling. We extend this picture and show that in the adiabatic limit of condensed-phase electron transfer, the onset of barrierless transition occurs at a lower driving force than predicted by the non-adiabatic Marcus formulation. Furthermore, in the adiabatic limit of the Marcus-inverted region, the standard mechanism of electron transfer becomes topologically forbidden. This behavior arises from a topological change in the mapping between the adiabatic and diabatic electronic surfaces, emerging precisely at the onset of the Marcus-inverted region. In this case, alternative mechanisms such as tunneling and non-radiative decay may dominate the rate, typically orders of magnitude slower than the rate calculated from Marcus theory.